OpenStudy (anonymous):

Can someone please help me with some integral homework before I shoot the crap out of my calculus book pretty please?! Equation in the forum

OpenStudy (anonymous):

\[\int\limits_{}^{}\ln (\sqrt[3]{x})\] That is x to the 3 root

OpenStudy (anonymous):

hi there

OpenStudy (anonymous):

hello = )

jimthompson5910 (jim_thompson5910):

it might help to think of the cube root of x as x^(1/3)

jimthompson5910 (jim_thompson5910):

then use the power rule

OpenStudy (anonymous):

ok do u want me to solve it or should i tell u how to do it

OpenStudy (anonymous):

ok I got as far to as thinking of x^(1/3) lol

jimthompson5910 (jim_thompson5910):

now use the rule that integral of x^n = (1/(n+1))*x^(n+1)

OpenStudy (anonymous):

if u solve it i give u medal and if i do so you give me one

OpenStudy (anonymous):

so do u mean put 1/3 in front of the ln of stuff and then put it in front of the entire integral? bc that's what I did and it said it was wrong = (

jimthompson5910 (jim_thompson5910):

there's no ln involved here

OpenStudy (anonymous):

u=x^1/3

OpenStudy (anonymous):

ok maybe you should write the whole thing out cuz now you're just confusing me, please, thank you

jimthompson5910 (jim_thompson5910):

does this rule look familiar